OpenStudy (anonymous):

What is the missing exponent? (6 )-4 = 620

OpenStudy (anonymous):

(6 )^-4 = 620 (the negative 4 is an exponent)

OpenStudy (campbell_st):

so are you looking for the exponent of 6...? as in \[(6^x)^{-4} = 620\]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

@campbell_st

OpenStudy (campbell_st):

ok... so using the power of a power law which means multiply the powers \[6^{-4x} = 620\] now you need logs to solve it take the log of both sides \[\ln(6^{-4x}) = \ln(620)\] applying the log laws for powers \[-4x \times \ln(6) = \ln(620)\] so \[-4x = \frac{\ln(620)}{\ln(6)}\] now solve for x

OpenStudy (anonymous):

I'm still confused, how do I solve for x?

OpenStudy (anonymous):

@campbell_st

OpenStudy (campbell_st):

so when you get a value for the log division, just divide it by -4